OpenStudy (anonymous):

If I get a 74.4 overall average and a test is worth 20 percent. There are 20 questions each worth 10 points. if failed how far will that drop overall score?

OpenStudy (amistre64):

74.4 = 80% of your grade ?? g/100 = 1% g/100 * 80 = 74.4 g = 7440/80

OpenStudy (amistre64):

i dont htink im reading the information correctly ....

OpenStudy (anonymous):

The test provides 200 total points. Since it makes up 20% of the score, and assuming scores aren't scaled down, that means the total number of points \(x\) for the class must be \(.2x=200~~\Rightarrow~~x=1000\). If \(y\) is the number of points you have now, and \(z\) is your score on the test, then the total class average will be \[Average=\frac{y+z}{1000}\] The current average, not counting the test points, is given to be 74.4% = 0.744, which means your current point total is \[.744=\frac{y}{800}~~\Rightarrow~~y=595.2\] What you have to do is find the average given some failing score on the test \(z\): \[Average=\frac{595.2+z}{1000}\] where \(0\le z\le200\). You probably have to use some known index for a "failing" score, maybe like a 60% on the test?